### v1

# On Geometric Evolution and Microlocal Regularity of the Navier–Stokes Equations

Abstract
We develop a geometric and microlocal framework for the Navier–Stokes equations by lifting the dynamics to the cosphere bundle of a Riemannian manifold. In this formulation, the velocity field and vorticity are represented as microlocal distributions whose evolution is governed by a transport–dissipation system generated by a canonical dynamical vector field described with differential forms. We introduce microlocal amplitudes, directional energy functionals, and monotone volume invariants on the compact phase space, which quantify directional concentration and alignment mechanisms associated with potential loss of regularity. The viscous term induces effective geometric diffusion on the cosphere bundle, yielding closed differential inequalities in a geometric setting. To capture the interaction between fluid deformation and geometry, we define an effective connection and curvature tensor encoding the influence of the symmetric velocity gradient. This structure gives rise to a Ricci-type geometric evolution at the microlocal level, controlled by a dissipative functional that excludes extreme directional concentration and persistent transverse stretching. Our results do not resolve the global regularity problem in full generality, but provide a coherent geometric mechanism that severely restricts blow-up scenarios compatible with viscous Navier–Stokes dynamics, reformulating the analytic regularity question as a problem of dissipative stability on a compact phase space.
Keywords
Navier–Stokes equations, microlocal analysis, cosphere bundle, geometric regularity criteria, effective Ricci curvature, dissipative dynamical systems, differential forms.
preprint: APS/123-QED
I Introduction: The Regularity Problem and Geometric Flows
The question about the existence and uniqueness of the solution space for the Navier-Stokes equations is one of the most significant and open research topics in mathematics and theoretical physics [11, 40], with a vast range of applications. Geometric and covariant formulations on Riemannian manifolds provide a coherent framework independent of coordinates [7, 13]. This clearly explains differential operators, the energy structure, the role of curvature, and the symmetries of the system through Killing fields [6]. This is because Hamiltonian formulations and differential forms, using weighted and covariant tensors with exterior derivatives [30], ensure the method works naturally, making diffusion and boundary terms appear more clearly in components with non-zero curvature.
For cases of two spatial dimensions with certain geometries and specific boundary conditions, existence, uniqueness, and continuity are well understood through frequency decomposition and invariant methods [14]. On compact differential manifolds with boundaries, there are strong and weak solutions [39]. Both are critical and converge exponentially to an equilibrium state given a certain evolution regime of the system, matching the Killing vector fields. Consequently, they are associated with invariant quantities that describe asymptotes, providing a model of the solution space near the origin. Similarly, the definition in spectral submanifolds is established in compact domains of the flow [5], with results showing the continuity of these invariants associated with the original differential manifold [36, 32].
In contrast, for methods with three or more spatial dimensions in critical spaces, there is strong evidence of non-uniqueness and geometric indefinitions in the solution space due to vorticities [2, 10]. More specifically, if the vorticity directions stay within a double cone in high-vorticity regions, the solutions remain regular; near singularities, the vorticity directions must spread across the unit sphere [31]. The quantitative classification of potential singularities and Carleman-type inequalities [28] provide testable lower and upper bounds near candidate blow-up times [8]. Classic explicit constructions show a blow-up in finite time in special stagnation-point forms (Euler and Navier-Stokes), driven by unlimited vorticity and infinite domains [9]. As a result, Fourier analysis and microlocal tools adapted to differential manifolds [27, 29] support the proofs of global regularity, geometric trapping [4], and hybrid discretizations in 3D domains [1]. These are based on eigenfunction and multilinear estimates adapted to curved spaces, and are considered methods to provide regularity and restrictions on possible singularities of the system [41, 26].
In this work, we reformulate the incompressible Navier–Stokes equations as a dissipative dynamical system defined globally on the cosphere bundle $S^{*}M$. Our goal is to classify blow-up mechanisms and obtain geometric criteria for conditional regularity using variational and Fourier methods. In this way, we find inequalities and invariants to show how the physical solution is reconstructed from the dynamics in $S^{*}M$.
II Covariant Formulation on Manifolds
To analyze the interaction between fluid dynamics and geometry, it is necessary to reformulate the Navier-Stokes equations in a fully covariant way on an oriented Riemannian differential manifold $(M,g)$. This manifold can be compact without a boundary or have a boundary with appropriate physical conditions, such that its curvatures and derivatives are bounded and the manifold is uniformly injective, equipped with a metric tensor. We denote its tangent and cotangent bundles as $TM$ and $T^{*}M$, respectively. In this section, we leave behind the traditional notation of vector calculus in favor of differential forms and the Hodge Laplacian operator, $\Delta_{H}$, which naturally adapts to the curvature of the space [13, 30]. This reformulation allows us to express momentum conservation and the incompressibility condition as geometric constraints that depend explicitly on the metric tensor. Therefore, the fluid no longer moves through a passive background but instead interacts with a dynamic geometric structure. Usually, for an incompressible fluid, we have the following system of equations using the covariant derivative $\nabla$ given a certain arbitrary connection $\Gamma$ on the Riemannian manifold $(M,g)$ induced by the metric [7].
$\left\{\begin{array}[]{cc}\partial_{t}u+\nabla_{u}u+\nabla p-\nu\Delta u-f=0\\ div(u)=0\end{array}\right$. (1)
Now, let’s consider the 1-forms associated with the velocity field and the external forces acting on the system, denoted as $\phi=\phi^{k}dx_{k}=u^{k}dx_{k}$ and $F=F^{k}dx_{k}$ respectively, with indices such that $k=1,2,\dots,\dim(M,g)$. This allows us to rewrite the original system (1) as follows
$\left\{\begin{array}[]{cc}\partial_{t}\phi+\mathcal{L}_{u}\phi+dP-\nu\Delta_{H}\phi-F=0\\ \delta\phi=0\end{array}\right$. (2)
where
  • $\mathcal{L}_{u}\phi=\iota_{u}d\phi+d(\iota_{u}\phi)=u^{j}(\nabla_{j}\phi_{i})+(\nabla_{i}u^{j})\phi_{j}$ is the Lie derivative of $\phi$ along the direction of $u$.
  • $\Delta_{H}=d\delta+\delta d$ is the Hodge Laplacian, which holds the Ricci tensor and the metric tensor within its definition.
  • $\delta:\Omega^{h}(M)\longmapsto\Omega^{h-1}(M)$ is the Hodge delta. It acts as the adjoint operator to the exterior derivative, such that $\delta=(-1)^{n(p+1)+1}\star d\ \star$ for a differential p-form. When applied to $\phi$, it results in $\delta\phi=-\nabla^{k}\phi_{k}=-g^{kj}\nabla_{j}\phi_{j}$.
Given these definitions and the system of equations we are looking at, if we use the incompressibility equation $\delta\phi=0$, the term for the Lie derivative regroups like this
$\mathcal{L}_{u}\phi=\nabla_{u}u^{k}+\frac{1}{2}d\|u\|^{2}_{L^{2}}$ (3)
Since the norm of $u$ appears here, it is helpful for establishing convergence criteria. As for the Hodge Laplacian, in local coordinates, we can write it as
$\Delta_{H}=(\nabla^{\star}\nabla\phi)_{k}+R^{k}_{ikj}g^{jk}\phi_{k}=\nabla^{i}\nabla_{i}\phi_{k}+R^{k}_{ikj}g^{jk}\phi_{k}$ (4)
By substituting this into the Navier-Stokes equations, we get
$\left\{\begin{array}[]{cc}\partial_{t}\phi+d(\frac{1}{2}\|u\|^{2}_{L^{2}}+P)-\nu(\nabla^{i}\nabla_{i}\phi_{k}+R^{k}_{ikj}g^{jk}\phi_{k})-F=0\\ g^{ik}\nabla_{i}\phi_{i}=0\end{array}\right$. (5)
This formulation makes the system of equations topologically more natural and clearer to analyze. Because of this, we define the covariant gradient of the velocity field
$A:=A_{ik}(dx^{i}\otimes dk^{k})=\nabla_{i}u_{k}(dx^{i}\otimes dk^{k})\in\Gamma(T^{\star}M\otimes TM)$ (6)
as a canonical decomposition induced by $g_{ik}$
$A=S+\Omega$ (7)
where
  • $S\in\Gamma(\text{Sym}^{2}T^{\star}M)$, such that $S_{ik}=\tfrac{1}{2}(\nabla_{i}u_{k}+\nabla_{k}u_{i})$, is a symmetric object by definition, representing the deformation tensor.
  • $\Omega\in\Omega^{2}(M)$ corresponds to the 2-form associated with the fluid’s rotation, defined as $\Omega_{ik}=\tfrac{1}{2}(\nabla_{i}u_{k}-\nabla_{k}u_{i})$.
The antisymmetric part $\Omega$ induces a differential 2-form which coincides, up to a constant, with the exterior derivative of the velocity 1-form. Indeed, by defining
$\omega:=d\phi\in\Omega^{2}(M)$ (8)
we get locally
$\omega_{ik}=\nabla_{i}u_{k}-\nabla_{k}u_{i}=2\Omega_{ik}$ (9)
To explicitly verify the relationship between the antisymmetric part of the gradient and the vorticity, consider the velocity 1-form
$\phi=u_{i}dx^{i}$ (10)
By the definition of the exterior derivative
$d\phi=\partial_{i}u_{j}\,dx^{i}\wedge dx^{j}=\tfrac{1}{2}(\partial_{i}u_{k}-\partial_{k}u_{i})\,dx^{i}\wedge dx^{k}$ (11)
By replacing partial derivatives with covariant derivatives, the connection terms cancel out due to antisymmetry, yielding
$d\phi=\tfrac{1}{2}(\nabla_{i}u_{k}-\nabla_{k}u_{i})\,dx^{i}\wedge dx^{k}$ (12)
Comparing this with the definition of $\Omega$, we conclude that,
$\omega:=d\phi=2\Omega\in\Omega^{2}(M)$ (13)
In arbitrary dimensions, the vorticity is a differential 2-form whose Hodge dual is
$\star\omega\in\Omega^{dim(M)-2}(M)$ (14)
Therefore, in three dimensions, the vorticity can be identified with a 1-form through the Hodge operator
$\xi:=\star\omega\in\Omega^{1}(M)$ (15)
which corresponds to the classical vorticity field. Consequently, the vorticity coincides with the dual of the antisymmetric part of the velocity gradient, confirming that the vortex stretching is fully encoded within the tensor $\Omega$.
Thus, the evolution of the vorticity field can be derived directly from the covariant formulation of the Navier–Stokes equations. By applying the exterior derivative to the momentum equation expressed in terms of the velocity 1-form $\phi$, and utilizing the identity $d^{2}\phi=0$, the evolution equation for the geometric vorticity is obtained [30].
$\partial_{t}\omega+\mathcal{L}_{u}\omega=\nu\Delta_{H}\omega+dF$ (16)
In the absence of external forces, or for conservative forces, the last term vanishes, and the equation adopts a purely geometric form.
In three dimensions, by applying the Hodge operator, this equation translates into an evolution equation for the classical vorticity field $\xi$
$\partial_{t}\xi+\mathcal{L}_{u}\xi=\nu\Delta\xi+(\nabla u)\cdot\xi$ (17)
where the last term represents the vortex stretching, induced exclusively by the symmetric part of the velocity gradient. To describe the dissipative structure of this dynamics, we introduce the vortex energy functional
$\mathcal{E}_{\omega}[\phi]:=\frac{1}{2}\int_{M}\|\omega\|^{2}_{L^{2}}dV=\frac{1}{2}\int_{M}\|d\phi\|^{2}_{L^{2}}dV$ (18)
This functional measures the total intensity of the fluid’s rotation and is invariant under gauge transformations $\phi\mapsto\phi+d\psi$ compatible with the incompressibility condition. Let $\phi_{\varepsilon}=\phi+\varepsilon\psi$, with $\psi\in\Omega^{1}(M)$ being smooth such that $\delta\psi=0$. Then,
$\frac{d}{d\varepsilon}\mathcal{E}_{\omega}[\phi_{\varepsilon}]\Big|_{\varepsilon=0}=\int_{M}\langle d\phi,d\psi\rangle\,dV=\int_{M}\langle\delta d\phi,\psi\rangle\,dV$ (19)
Therefore, the variational gradient of the functional is given by
$\frac{\delta\mathcal{E}_{\omega}}{\delta\phi}=\delta d\phi=\Delta_{H}\phi$ (20)
Consequently, the viscous term of the Navier–Stokes equations can be interpreted as the gradient descent flow of the vortex energy functional with respect to the $L^{2}$ metric induced by $g$. In this sense, the dynamics of an incompressible fluid is presented as a dissipative system whose regularization mechanism is governed by a geometric variational principle. By differentiating with respect to time and using the evolution equation for $\omega$, we obtain the dissipation identity
$\frac{d}{dt}\mathcal{E}_{\omega}(t)=-\nu\int_{M}\|\nabla\omega\|^{2}_{L^{2}}dV\leq\gamma\Big|_{\gamma=0}=0$ (21)
which shows that viscosity imposes a strict monotonicity on the vortex energy, suggesting that the regularity of the system is intimately linked to the existence of dissipative geometric quantities that control the complexity of the flow.
Although the covariant formulation clarifies the geometric structure of the system, the regularity problem remains open in three dimensions. In particular, the vorticity equation reveals that potential amplification is associated with directional stretching, suggesting that the analysis must incorporate not only the magnitude of the vorticity but also its directional structure.
This motivates lifting the problem to a space where cotangent directions can be treated geometrically in a global manner, leading us to the cosphere bundle $S^{*}M$ through the cotangent bundle [20].
III Microlocal Lifting and the Cosphere Bundle
As discussed in the previous section, the natural space for this lift is the cotangent bundle $T^{*}M$, equipped with its canonical symplectic structure given by a projection. Let
$\pi:T^{*}M\longmapsto M,\qquad\pi(x,\xi)=x$ (22)
denote the natural projection, where each fiber $T_{x}^{*}M$ represents the space of co-vectors at the point $x\in M$.
Since the system is invariant under homogeneous rescaling of the cotangent variables, it is convenient to restrict our attention to the co-sphere bundle, defined as
$S^{*}M:=\{(x,\xi)\in T^{*}M:\|\xi\|_{g}=1\}$ (23)
which can be interpreted as the space of unit cotangent directions over $M$. There exists an isomorphism that does not change the homotopy class of the differential manifold considered:
$S^{*}M\cong M\times S^{dim(M)-1}$ (24)
From this perspective, a point in $S^{*}M$ simultaneously encodes a position in physical space and a privileged direction in frequency space. Therefore, the cotangent bundle $T^{*}M$ possesses a canonical symplectic structure given by the 2-form
$\Lambda=d\lambda$ (25)
where $\lambda=\xi_{i}\,dx^{i}$ is a 1-form. This structure is independent of the metric and encodes the microlocal geometry of the system, allowing for the description of the propagation of directional structures, such as vorticity, in a global and geometrically coherent manner.
Consequently, since fluid dynamics is implicitly related to geometry, it is logical to argue that the fluid couples to the metric or forms part of it, shaping the evolution of the manifold. Thus, a natural decomposition of the tangent bundle of $T^{*}M$ into horizontal and vertical subspaces is induced:
$TT^{*}M\simeq H\oplus V$ (26)
where the horizontal distribution is determined by the corresponding connection and the vertical distribution corresponds to purely cotangent variations.
This decomposition allows for the definition of coframes adapted to the microlocal lift and provides a natural geometric basis for lifting differential operators defined on $M$ to the cosphere bundle.
The vertical distribution $V$ is given by
${V}_{(x,\xi)}=\ker(d\pi)_{(x,\xi)}\simeq T_{\xi}(T_{x}^{*}M)$ (27)
While for the horizontal distribution, the connection $\Gamma$ defines it such that $H\subset TT^{*}M$ is characterized by the property that for every $X_{x}\in T_{x}M$, there exists a unique vector $X^{H}_{(x,\xi)}\in T_{(x,\xi)}T^{*}M$ such that
$d\pi\big[X^{H}_{(x,\xi)}\big]=X_{x},\qquad\Gamma_{X_{x}}\xi=0$ (28)
In this way, the decomposition is invariant under isometric diffeomorphisms and provides a natural geometric basis for lifting operators defined on $M$ to $T^{*}M$ and $S^{*}M$. To prove this, let $\varphi:M\longmapsto M$ be an isometric diffeomorphism, that is,
$\varphi^{*}g=g$ (29)
Then the horizontal distribution $H\subset TT^{*}M$ is invariant under the induced diffeomorphism $\tilde{\varphi}$, which implies that
$d\tilde{\varphi}\big[H_{(x,\xi)}\big]=H_{\tilde{\varphi}(x,\xi)}$ (30)
Given $X^{H}_{(x,\xi)}\in H_{(x,\xi)}$, we have $d\pi(X^{H})=X_{x}$ and $\Gamma_{X_{x}}\xi=0$. Since $\varphi$ is an isometry, it preserves the connection, so
$\Gamma_{d\varphi(X_{x})}\big[(d\varphi^{-1})^{*}\xi\big]=0$ (31)
Furthermore,
$d\pi\big[d\tilde{\varphi}(X^{H})\big]=d\varphi(X_{x})$ (32)
By the uniqueness of the horizontal lift, it follows that $d\tilde{\varphi}(X^{H})\in H_{\tilde{\varphi}(x,\xi)}$. Thus, the invariance under isometric diffeomorphisms is proven, as originally intended. Following this, the velocity field $u\in\Gamma(TM)$ canonically induces a dynamical field on $T^{*}M$, defined by
$X_{u}(x,\xi)=u^{i}(x)\,X_{i}^{H}-(\nabla_{k}u^{i})(x)\,\xi_{i}\,\frac{\partial}{\partial\xi_{k}}$ (33)
which simultaneously describes spatial transport and the directional evolution of covectors under the action of the flow [29].
From a geometric perspective, the dynamical vector field $X_{u}$ endows the cotangent bundle with a structure closely analogous to the configuration space of an unquantized field theory [34, 38]. In this interpretation, the microlocal variables $(x,\xi)\in T^{*}M$ play the role of generalized phase-space coordinates, while the flow generated by $X_{u}$ encodes both spatial transport and internal directional evolution.
This viewpoint naturally leads to the introduction of an effective connection adapted to the flow. In particular, the velocity field $u$ induces a deformation of the connection through
$(\Gamma_{u})^{j}_{ik}:=\Gamma^{j}_{ik}+g^{\prime}_{ik}\,u^{j}$ (34)
which incorporates the influence of the fluid transport into the geometric structure of the bundle [3]. Associated with this deformation, we consider an effective metric of the form
$g^{\prime}_{ik}=g_{ik}+\alpha\,\int_{t_{0}}^{t}S_{ik}(s)\ ds,\qquad S_{ik}=\frac{1}{2}\bigl(\nabla_{i}u_{k}+\nabla_{k}u_{i}\bigr)$ (35)
where $\alpha$ is a fixed geometric coupling parameter [21]. This metric deformation encodes the interaction between the background geometry and the fluid deformation tensor.
Under this construction, the curvature tensor associated with the effective connection $\Gamma^{u}$ is given in local coordinates by
$(R^{u})^{l}{}_{ijk}$ $=\partial_{i}(\Gamma^{l}_{jk}+g^{\prime}_{jk}u^{l})-\partial_{j}(\Gamma^{l}_{ik}+g^{\prime}_{ik}u^{l})$
$\quad+(\Gamma^{m}_{jk}+g^{\prime}_{jk}u^{m})(\Gamma^{l}_{im}+g^{\prime}_{im}u^{l})$
$\quad-(\Gamma^{m}_{ik}+g^{\prime}_{ik}u^{m})(\Gamma^{l}_{jm}+g^{\prime}_{jm}u^{l}).$ (36)
This decomposes as
$(R^{u})^{l}{}_{ijk}=R^{l}{}_{ijk}+\mathcal{A}^{l}{}_{ijk}+\mathcal{Q}^{l}{}_{ijk}$, (37)
where $R^{l}{}_{ijk}$ is the Riemann curvature of $g$
$\mathcal{A}^{l}{}_{ijk}=(\nabla_{i}g^{\prime}_{jk}-\nabla_{j}g^{\prime}_{ik})u^{l}+g^{\prime}_{jk}\nabla_{i}u^{l}-g^{\prime}_{ik}\nabla_{j}u^{l}$ (38)
and $\mathcal{Q}^{l}{}_{ijk}$ denotes quadratic terms in $u$,
$\mathcal{Q}^{l}{}_{ijk}=g^{\prime}_{jk}g^{\prime}_{im}u^{m}u^{l}-g^{\prime}_{ik}g^{\prime}_{jm}u^{m}u^{l}$ (39)
The effective Ricci tensor is obtained by contraction,
$R^{u}_{jk}=(R^{u})^{i}{}_{jik}$ (40)
Explicitly,
$R^{u}_{jk}$ $=R_{jk}+\nabla^{i}g^{\prime}_{jk}\,u_{i}-\nabla^{i}g^{\prime}_{ik}\,u_{j}+g^{\prime}_{jk}\nabla^{i}u_{i}-g^{\prime}_{ik}\nabla^{i}u_{j}$
$\quad+g^{\prime}_{jk}g^{\prime}_{im}u^{i}u^{m}-g^{\prime}_{ik}g^{\prime}_{jm}u^{i}u^{m}$ (41)
Under the incompressibility condition $\nabla^{i}u_{i}=0$ and at leading order in the metric deformation, this reduces to
$R^{u}_{jk}\simeq R_{jk}+\alpha\,\nabla^{i}\bigl(\nabla_{j}S_{ik}+\nabla_{k}S_{ij}-\nabla_{i}S_{jk}\bigr).$ (42)
making explicit the coupling between curvature and deformation. In terms of the effective Ricci tensor $R^{u}_{ik}$, the metric evolution over time is
$\partial_{t}g^{\prime}_{i}k=\partial_{t}g_{ik}+\alpha S_{ik}$ (43)
may be formally rewritten as a Ricci-type flow,
$\partial_{t}g^{\prime}_{ik}+2R^{u}_{ik}=\mathcal{E}_{ik}$ (44)
where $\mathcal{E}_{ik}$ represents higher-order and transport-induced correction terms controlled by the deformation tensor $S$. Then, by definition $\mathcal{E}_{ik}$ may be regarded as
$\mathcal{E}_{ik}=\partial_{t}g_{ik}+\alpha S_{ik}+2R^{u}_{ik}$ (45)
This comparison does not assert that the physical metric evolves by Ricci flow, but rather that the Navier–Stokes dynamics induces a Ricci-type geometric structure at the microlocal level, where viscous dissipation enforces curvature control and prevents the formation of singular geometric configurations compatible with blow-up [4]. Then, the correction tensor $\mathcal{E}_{ik}$ admits the canonical decomposition
$\mathcal{E}_{ik}=\mathcal{E}^{(I)}_{ik}+\mathcal{E}^{(II)}_{ik}+\mathcal{E}^{(III)}_{ik}+\mathcal{E}^{(IV)}_{ik}$ (46)
where
$\mathcal{E}^{(I)}_{ik}$ $=\partial_{t}g_{ik}+2R_{ik}$ (47)
$\mathcal{E}^{(II)}_{ik}$ $=\alpha S_{ik}$ (48)
$\mathcal{E}^{(III)}_{ik}$ $=2\alpha\nabla^{j}(\nabla_{i}S_{jk}+\nabla_{k}S_{ji}-\nabla_{j}S_{ik})$ (49)
$\mathcal{E}^{(IV)}_{ik}$ $=2\mathcal{Q}_{ik}(u)$ (50)
Therefore, the tensor $\mathcal{E}_{ik}$ measures the deviation of the effective metric evolution from a pure Ricci flow, encoding transport, deformation, and higher-order nonlinear effects induced by the fluid dynamics. The first term of $\mathcal{E}_{ik}$ goes for geometric background, the second one for instantaneous deformation, the third one for geometric diffusion and the fourth one for non-linear transport.
IV Geometric Evolution Operators and Microlocal Representation of Solutions
By homogeneity, the field $X_{u}(x,\xi)$ is tangent to $S^{*}M$ and therefore defines a well-defined dynamics on the cosphere bundle, where $\|\xi\|_{g^{\prime}}$ is preserved. This will be useful because, at this point, it is beneficial to perform a Fourier transform on the system [5, 6]. Since the classical Fourier transform is not globally defined on compact manifolds, we employ a microlocal transform defined locally in normal coordinates and extended globally as a function on $T^{*}M$. For a 1-form $\phi\in\Omega^{1}(M)$, we define
$\tilde{\phi}(x,\xi)=\int_{T_{x}M}e^{-i\langle\xi,y\rangle}\phi(\exp_{x}y)\,\chi(y)\,dy$ (51)
where $\chi$ is a smooth cutoff function. Restricting by homogeneity to $S^{*}M$, we obtain $\tilde{\phi}\in C^{\infty}(S^{*}M)$.
Applying the microlocal transform to the incompressible covariant system and projecting onto transverse directions to eliminate the pressure term, we obtain the microlocal equation
$\partial_{t}\tilde{\phi}+\mathcal{L}_{X_{u}}\tilde{\phi}+\nu\|\xi\|_{g}^{2}\,\tilde{\phi}=\tilde{R}+\tilde{F},\qquad\langle\tilde{\phi},\xi\rangle_{g}=0$ (52)
Equation (52) is a linear non-autonomous dissipative system on $S^{*}M$. Here, the term corresponding to the pressure $dP$ does not contribute to the system’s dynamics, acting as a Lagrange multiplier that imposes the incompressibility condition, existing only in the radial direction of the cotangent bundle. Therefore, we define the incompressible microlocal projector
$\Pi_{\xi}:T^{\star}_{x}M\longmapsto\xi^{\perp}\subset T^{\star}_{x}M$ (53)
given by
$\Pi_{\xi}(\eta)=\eta-\frac{\langle\eta,\xi\rangle_{g}}{\|\xi\|_{g}^{2}}\xi$ (54)
which yields the conditions
$\Pi_{\xi}(\xi)=0,\quad\Pi_{\xi}(dP)=0$ (55)
This causes the pressure term to vanish when restricting the equation to the cosphere bundle and projecting onto transverse directions. In fact, the condition of the divergence transform also justifies the defined projection and confirms that the dynamics evolve naturally in $S^{*}M$.
In order to solve the microlocal system on the cosphere bundle $S^{*}M$, we define the Green operator $G(t,s)$ as the fundamental solution of the non-autonomous dissipative equation [25, 38]. Let $Q(t)=\mathcal{L}_{X_{u}}+\nu\|\xi\|_{g}^{2}$ be the differential operator acting on sections of the bundle. The Green operator satisfies
$\left(\partial_{t}+\mathcal{L}_{X_{u}}+\nu\|\xi\|_{g}^{2}\right)G(t,s)=\delta(t-s)\mathbb{I}$ (56)
where $\mathbb{I}$ is the identity operator on $C^{\infty}(S^{*}M)$.
Let $\Psi_{t,s}:S^{*}M\longmapsto S^{*}M$ be the flow generated by the dynamical vector field $X_{u}$. This flow represents the lifting of the fluid’s trajectories to the phase space. For any initial microlocal distribution $\tilde{\phi}_{0}$, the pure transport part without dissipation is given by the pull-back $(\Psi_{t,s}^{*}\tilde{\phi}_{0})(x,\xi)=\tilde{\phi}_{0}(\Psi_{s,t}(x,\xi))$.
In the general case, let $\tilde{f}\in C^{\infty}(S^{*}M\times N)$ represent the total microlocal source term, encompassing both external forces and non-linear residuals with N{N} is a general manifold representing additional degrees of freedom, parameters or internal symmetries of the system. In the relativistic limit, $N$ is identified with the timelike component of the worldline, and the physical fields are recovered by the push-forward integrating over the fibers. Then, the Green operator is constructed by coupling the geometric transport with the exponential dissipation factor induced by the viscosity and the metric.
$G(t,s)[\tilde{f}](x,\xi)=\exp\left(-\int_{s}^{t}\nu\|\xi(\tau)\|_{g}^{2}\,d\tau\right)\tilde{f}(\Psi_{s,t}(x,\xi))$ (57)
where
  • $\xi(\tau)$ denotes the evolution of the covector along the integral curves of $X_{u}$ in the cotangent bundle.
  • $\tilde{f}$ is understood as a section whose values in N{N} are preserved along the flow $\Psi_{s,t}$ unless acted upon by internal dynamics, being part of the phase space for the system, describing the Fourier transform of the external forces and non-linear residuals of curvature.
To verify that $G(t,s)$ is the fundamental solution, we observe, at $t=s$, the flow $\Psi_{s,s}$ is the identity and the integral in the exponent vanishes, thus $G(s,s)=\mathbb{I}$. Applying the operator $\partial_{t}+\mathcal{L}_{X_{u}}$ to the transported term, we have by the property of the Lie derivative and the flow
$(\partial_{t}+\mathcal{L}_{X_{u}})\left[\tilde{f}(\Psi_{s,t}(x,\xi))\right]=0$ (58)
which follows from the fact that $\Psi_{s,t}$ is the solution to the transport equation associated with $X_{u}$. Differentiating the exponential factor with respect to $t$ yields the term $\nu\|\xi\|_{g}^{2}$. Summing these contributions, we obtain
$(\partial_{t}+\mathcal{L}_{X_{u}}+\nu\|\xi\|_{g}^{2})G(t,s)\\ =(-\nu\|\xi\|_{g}^{2}+\nu\|\xi\|_{g}^{2}+\mathcal{L}_{X_{u}}-\mathcal{L}_{X_{u}})G(t,s)=0$ (59)
for $t>s$ and it converges for the delta distribution at $t=s$. The operator thus acts as a low-pass filter in the phase space, where the dissipation follows the geometric trajectories.
Thus, the solution to the transformed microlocal system can be formally written by solving for $\tilde{\phi}$ over the time variable as
$\tilde{\phi}(t)=\mathcal{T}\bigg[\exp\left(-\int_{t_{0}}^{t}\big(\mathcal{L}_{X_{u}}+\nu\|\xi\|_{g}^{2}\big)\,ds\right)\tilde{\phi}_{0}\bigg]\\ +\int_{t_{0}}^{t}G(t,s)\big[\tilde{R}(s)+\tilde{F}(s)\big]\,ds$ (60)
where
  • $\mathcal{T}$ is the time-ordering operator.
  • $G(t,s)$ is the associated Green’s operator.
  • $\tilde{\phi}_{0}$ is the initial microlocal data.
Finally, the physical solution is recovered by integrating over the cotangent directions
$\phi(x,t)=\int_{S_{x}^{\star}M}\tilde{\phi}(x,t,\xi)\,d\xi$ (61)
which defines a geometrically canonical reconstruction of the velocity field.
This formulation allows for the study of the system’s regularity through the microlocal behavior of $\tilde{\phi}$, reducing the global analysis of Navier-Stokes to the dynamical properties of the flow induced on $S^{*}M$. Analogously, we follow the same procedure for vorticity, obtaining the following equation
$\partial_{t}\tilde{\omega}+\mathcal{L}_{X_{u}}\tilde{\omega}=-\nu\|\xi\|_{g}^{2}\tilde{\omega}+\tilde{R}_{\omega}+i\xi\wedge\tilde{F},\qquad\xi\wedge\tilde{\omega}=0$ (62)
The condition $\xi\wedge\tilde{\omega}=0$ is the microlocal version of $d\omega=0$. Equation (62) is linear and dissipative, and its solution is formally written as
$\tilde{\omega}(t)=\mathcal{T}\bigg[\exp\left[-\int_{t_{0}}^{t}(\mathcal{L}_{X_{u}}+\nu\|\xi\|_{g}^{2})\,ds\right]\tilde{\omega}_{0}\bigg]\\ +\int_{t_{0}}^{t}G(t,s)\big[\tilde{R}\omega(s)+i\xi\wedge\tilde{F}(s)\big]ds$ (63)
If $\tilde{\phi}(t)$ is a solution to (60), then $\tilde{\omega}(t):=d\tilde{\phi}(t)$ satisfies the microlocal vorticity equation (62). This is because the exterior derivative commutes with the differential operators used in the system for smooth differential forms, and $d(dP)=0$; therefore, applying the exterior derivative to (52) yields exactly (62) with a solution reconstructed also from
$\omega(x,t)=\int_{S_{x}^{*}M}\tilde{\omega}(x,t,\xi)\,d\xi$ (64)
It is inferred that the microlocal lift of the Navier-Stokes equations produces a well-defined dissipative dynamical system on the cosphere bundle, whose solution admits an explicit representation through evolution operators. The exact compatibility between the velocity and vorticity equations guaranties the consistency of the formalism. This approach reduces the global regularity problem to the control and convergence of geometric microlocal invariants, the analysis of which is addressed in the following section.
V Stability, Energy and Geometric Criteria for Regularity and Smoothness
In the previous sections, the incompressible Navier–Stokes equations were reformulated as a linear, non-autonomous, dissipative dynamical system acting on microlocal lifts of differential forms over the compact phase space $S^{*}M$. We now exploit this structure to derive quantitative stability estimates and geometric criteria for regularity [28]. The central idea of this section is that any potential singular behavior of the physical solution must manifest itself as a geometric instability of the microlocal dynamics on $S^{*}M$ [29, 7]. Viscosity, compactness, and the geometric nature of the lift combine to impose strong restrictions on such instabilities.
Throughout this section, $\lambda\in\{\phi,\omega\}$ denotes either the velocity $1$-form or the vorticity $2$-form on $M$, and $\tilde{\lambda}$ its microlocal lift to $S^{*}M$. All norms and contractions are taken with respect to the effective metric $g^{\prime}$ introduced earlier. Since $S^{*}M$ is compact and endowed with a canonical contact structure, all integrals below are well-defined without further assumptions.
We begin by defining the fundamental local geometric quantity governing the analysis. The microlocal amplitude of $\tilde{\lambda}$ is defined pointwise by the contraction
$|\tilde{\lambda}|_{g^{\prime}}^{2}:=\langle\tilde{\lambda},\tilde{\lambda}\rangle_{g^{\prime}}=\iota_{\tilde{\lambda}^{\sharp}}\tilde{\lambda}$, (65)
where $\tilde{\lambda}^{\sharp}$ denotes the metric dual with respect to $g^{\prime}$. This quantity measures the intensity of the lifted differential form along a given cotangent direction. Unlike classical norms, it retains full directional information encoded in the cosphere bundle.
The corresponding global geometric functional is obtained by pairing this density with the canonical Liouville volume form on $S^{*}M$,
$\mu_{S^{*}M}=\alpha\wedge(d\alpha)^{n-1}$ (66)
where $\alpha$ is the contact form induced from the Liouville $1$-form on $T^{*}M$. We define the microlocal energy functional
$\mathcal{V}_{\lambda}(t):=\int_{S^{*}M}|\tilde{\lambda}(x,\xi,t)|_{g^{\prime}}\,\mu_{S^{*}M}$ (67)
Geometrically, $\mathcal{V}_{\lambda}$ measures the total directional spread of the field $\lambda$ in phase space. Physically, it quantifies how much the solution concentrates along preferred cotangent directions, independently of its spatial localization.
We now derive a balance law for $\mathcal{V}_{\lambda}$. The microlocal evolution equation satisfied by $\tilde{\lambda}$ is
$\partial_{t}\tilde{\lambda}+\mathcal{L}_{X_{u}}\tilde{\lambda}+\nu\|\xi\|_{g}^{2}\,\tilde{\lambda}=\tilde{R}[\tilde{\lambda}]+\tilde{C}[\tilde{\lambda}]+\tilde{F}$ (68)
where $X_{u}$ is the canonical lifted transport field on $S^{*}M$. We take the pointwise inner product with $\tilde{\lambda}$ and use bilinearity to obtain
$\frac{1}{2}\partial_{t}|\tilde{\lambda}|_{g^{\prime}}^{2}+\frac{1}{2}\mathcal{L}_{X_{u}}|\tilde{\lambda}|_{g^{\prime}}^{2}+\nu\|\xi\|_{g}^{2}|\tilde{\lambda}|_{g^{\prime}}^{2}\\ =\langle\tilde{R}[\tilde{\lambda}],\tilde{\lambda}\rangle+\langle\tilde{C}[\tilde{\lambda}],\tilde{\lambda}\rangle+\langle\tilde{F},\tilde{\lambda}\rangle$ (69)
Each term on the right-hand side admits a sharp geometric bound. Since the curvature term is zeroth order in $\xi$,
$|\langle\tilde{R}[\tilde{\lambda}],\tilde{\lambda}\rangle|\leq\|\mathrm{Ric}\|_{L^{\infty}(M)}\,\|\tilde{\lambda}\|_{g^{\prime}}^{2}$ (70)
The commutator term, arising from non-exact microlocalization, satisfies
$|\langle\tilde{C}[\tilde{\lambda}],\tilde{\lambda}\rangle|\leq C\,\|\nabla u\|_{L^{\infty}(M)}\,\|\tilde{\lambda}\|_{g^{\prime}}^{2}$ (71)
Finally, the forcing term is controlled linearly,
$|\langle\tilde{F},\tilde{\lambda}\rangle|\leq\|\tilde{F}\|\,\|\tilde{\lambda}\|_{g^{\prime}}$ (72)
We now integrate the balance identity over $S^{*}M$. Since the Liouville measure is invariant under the flow generated by $X_{u}$, that is,
$\mathcal{L}_{X_{u}}\mu_{S^{*}M}=0$ (73)
the transport term drops out after integration. We obtain
$\frac{d}{dt}\mathcal{V}_{\lambda}(t)+\nu\int_{S^{*}M}\|\xi\|_{g}^{2}\|\tilde{\lambda}\|_{g^{\prime}}^{2}\,\mu_{S^{*}M}\leq C\|\nabla u\|_{L^{\infty}}\mathcal{V}_{\lambda}(t)\\ +\|\tilde{F}\|_{L^{1}(S^{*}M)}$ (74)
This inequality expresses a geometric fact, microlocal energy cannot accumulate without being dissipated, unless the deformation of the flow becomes uncontrolled. In particular, in the absence of forcing, $\mathcal{V}_{\lambda}$ is monotone up to a deformation-controlled factor.
To make this mechanism more explicit, recall that the velocity gradient decomposes as $\nabla u=S+\Omega$, where only the symmetric part $S$ contributes to stretching. Restricting the evolution to integral curves of $X_{u}$ yields the pointwise inequality
$\frac{d}{dt}\|\tilde{\lambda}\|\leq\|S(t)\|_{L^{\infty}}\,\|\tilde{\lambda}\|-\nu\|\tilde{\lambda}\|$ (75)
with solution
$\|\tilde{\lambda}(t)\|\leq\|\tilde{\lambda}(0)\|\exp\!\left[\int_{t_{0}}^{t}(\|S(s)\|_{L^{\infty}}-\nu)\,ds\right]$ (76)
showing that any amplification of microlocal amplitude requires a persistent failure of temporal integrability of $\|S\|_{L^{\infty}}$.
We now specialize to the vorticity $\tilde{\omega}$, whose geometric structure allows for finer control. Since $\tilde{\omega}$ is a closed microlocal $2$-form, we introduce the functional
$W[\tilde{\omega}]=\int_{S^{*}M}\left(\tau\,|d^{\perp}\log|\tilde{\omega}||^{2}+\log|\tilde{\omega}|\right)\mu_{S^{*}M}$ (77)
where $d^{\perp}$ denotes the vertical exterior derivative induced by the horizontal–vertical splitting of $TS^{*}M$, and $\tau>0$ is a dimensional parameter. This functional penalizes sharp angular gradients of vorticity and quantifies directional irregularity.
A direct differentiation in time, combined with integration by parts along the fibers [30] and the compactness of $S^{*}M$, yields the dissipation inequality
$\frac{d}{dt}W[\tilde{\omega}]\leq\int_{S^{*}M}\left(\tau\,|d^{\perp}\log|\tilde{\omega}||^{2}-\tau\,\|S\|\,|\tilde{\omega}|-\nu\right)\mu_{S^{*}M}$ (78)
This inequality shows that viscosity and geometric diffusion suppress directional oscillations of vorticity, unless sustained stretching dominates dissipation.
Since $S^{*}M$ is compact, the above estimates imply uniform control of $\tilde{\omega}$ in $L^{\infty}(S^{*}M)$, and therefore of the reconstructed physical vorticity [4]. Consequently, any loss of regularity must be accompanied by a breakdown of one of the assumptions used above.
We arrive at a geometric classification of admissible blow-up mechanisms [8]. A singularity can occur only if at least one of the following phenomena takes place,
  • Extreme directional concentration of microlocal vorticity in $S^{*}M$.
  • Persistent loss of vortex alignment encoded in the symmetric deformation tensor.
  • Failure of temporal integrability of $\|\nabla u\|_{L^{\infty}}$.
Viscous dissipation excludes all other scenarios by enforcing monotonicity of microlocal geometric invariants.
In conclusion, the Navier-Stokes equations admit a rigid geometric reformulation in which regularity is equivalent to the stability of a dissipative dynamical system on a compact phase space. Blow-up is not a purely analytic accident, but a geometric instability that must overcome strong dissipative and structural constraints.
VI Discussion and Conclusions
The theoretical approach developed in this work addresses the Navier–Stokes regularity problem through a synthesis of microlocal analysis and Riemannian geometry. The central structural step is the lifting of the dissipative dynamics from the physical manifold $M$ to the cosphere bundle $S^{*}M$, which resolves several ambiguities inherent in formulations based solely on physical space. In contrast with classical regularity theories focused on the control of Sobolev norms and vorticity amplitudes [11, 40], the present framework emphasizes directional alignment and geometric dissipation as primary mechanisms governing regularity. This perspective is consistent with geometric trapping phenomena [4] and extends them to three dimensions through the use of the Hodge Laplacian and its variational structure [7].
The geometric role of curvature, revealed through the Weitzenböck identity and the coupling of the Ricci tensor to vorticity dynamics, shows that the topology of $M$ actively contributes to dissipation rather than serving as a passive background [3, 7]. The monotonicity of the vortex energy functional $\mathcal{E}_{\omega}$ and the associated microlocal inequalities restrict admissible blow–up scenarios in a manner compatible with known geometric constraints [8, 31]. In this setting, the vortex stretching term is reinterpreted not as an intrinsically destabilizing mechanism, but as a geometric deformation acting on a compact fiber, where viscous effects induce effective angular diffusion. The microlocal transport–dissipation structure is consistent with the analysis of lifted operators and unique continuation phenomena developed in [29], and aligns with energy–aware geometric decompositions of the Navier–Stokes system [6].
For that, this paper does not solve the Navier–Stokes regularity problem. Instead, it introduces a coherent and geometrically meaningful framework that recasts the question in microlocal and dynamical terms, where regularity is linked to directional alignment and dissipative stability on a compact phase space [29]. In particular, the formulation is compatible with recent results on non–uniqueness of smooth or weak solutions arising from critical data [10], this can be interpreted as manifestations of microlocal flexibility within the constraints imposed by energy dissipation and geometric monotonicity. The present framework does not preclude non–uniqueness at low regularity; rather, it provides a mechanism to distinguish dynamically stable microlocal configurations from those that are highly sensitive to perturbations and thus prone to branching or loss of uniqueness.
From this viewpoint, regularity is not characterized solely by uniqueness or smoothness in physical space, but by the persistence of controlled microlocal structure along the lifted flow. Singular or non–unique behaviors must then correspond to a breakdown of directional control or of dissipative balance in $S^{*}M$, rather than to uncontrolled growth of classical norms alone. This reformulation clarifies how geometric dissipation and microlocal transport can coexist with modern flexibility results, while still yielding meaningful regularity criteria for physically relevant solutions.
A complementary geometric interpretation is obtained by considering the lifted dynamics on the extended manifold $S^{*}M\times N$, where $N\subset\mathbb{R}$ denotes the temporal interval. In this formulation, the temporal boundary components of $N$ act as Cauchy hypersurfaces for the evaluation of microlocal energies and functionals [39]. By Stokes’ theorem, variations of the geometric invariants introduced in this work reduce to boundary fluxes [30], showing that changes in microlocal energy and directional structure are governed by boundary contributions rather than interior concentration mechanisms. Preservation of smoothness can therefore be interpreted as the absence of singular behavior in the asymptotic microlocal distribution, excluding concentration at infinity in phase space [31, 4].
Despite the structural coherence of the microlocal framework developed here, several important limitations and open directions remain. First, while the lifting to $S^{*}M$ provides refined control over directional alignment and dissipative mechanisms, it inevitably entails a loss of direct information when projecting back to physical space. Quantifying precisely which microlocal features are preserved under reconstruction, and which may be lost or blurred, remains an open problem. In particular, sharper criteria are needed to distinguish benign directional dispersion from genuinely unstable microlocal configurations.
Second, the present analysis is carried out on a fixed Riemannian manifold and does not address the possibility of topological or geometric transitions in the underlying space. Understanding how changes in topology, curvature, or bundle structure affect microlocal stability and dissipation represents a natural and largely unexplored direction. Such effects may be relevant in flows with time–dependent geometry or in singular limits where geometric control degenerates.
Finally, although the derived inequalities severely restrict admissible blow–up scenarios, they do not yet yield a complete classification of singularity formation mechanisms. Developing refined microlocal criteria that can discriminate between transient directional amplification and genuinely singular behavior, as well as identifying optimal strategies for detecting or preventing microlocal concentration, remain important challenges for future work. Addressing these questions may require combining the present geometric framework with finer second microlocalization techniques or with adaptive topological invariants capable of capturing changes in phase–space structure. Making this approach a geometric reorganization of the regularity problem rather than a final resolution [37]. Its structure may nevertheless open new avenues for understanding singularity formation and dissipative stability from a geometric perspective [8].
Beyond its theoretical implications, the microlocal framework suggests potential applications and development in structure–preserving numerical schemes and geometry–aware computational fluid dynamics, where directional instabilities are controlled through intrinsic geometric mechanisms rather than arbitrary regularization [1, 6, 15, 5]. These ideas may be useful in turbulence modeling, data-driven reduced-order models, and multi-scale simulations in cluster environments where know and preserve the appropriate geometric properties and its structure is essential for long-time accuracy.
VII AI Statement
The author confirms that every intellectual contribution, original mathematical and physical derivations, theorem, proofs, and physical hypotheses presented in this manuscript are the sole work of the author.
Generative Artificial Intelligence (AI) tools, specifically the Large Language Model ChatGPT, were used exclusively as digital research assistants for the following purposes: refining academic language and structuring the references into alphabetic order. The AI did not perform any original scientific calculations, nor proposed any scientific content as hypotheses, proofs or others. The author has reviewed, edited, and validated the accuracy of all output generated by the AI.
Acknowledgements.
The author is profoundly grateful to PhD. Pedro Díaz-Navarro because of his openess and creativity to talk. In the same way as Octavio Rojas-Quesada for every new point of view and Celicia Ledezma-Porras for her geometric insight some years ago. Moreover, family, some friends and every guide, person… who has stumble with me along the paths, thank you all.